Problems 23-25. Use the characteristics of the binomial experiment below to answer the questions. Commuter distance. Sixty-two percent of workers say they commute more than 15 miles to work. You randomly select 20 workers and ask if they commute more than 15 miles to work.

23. What is the probability that exactly ten workers say they commute more than 15 miles to work?

24. What is the probability that more than 16 workers say they commute more than 15 miles to work?

25. What is the probability that between eight and ten (inclusive) workers say they commute more than 15 miles to work?
Answer #1

This is a binomial distribution question with
n = 20
p = 0.62
q = 1 - p = 0.38
where

a) x = 10

P(X = 10) = 0.0974
b) x = 16
P(X > 16) = P(X = 17) + P(X = 18) + P(X = 19) + P(X = 20)


P(X > 16) = 0.0245
c) x1 = 8
x2 = 10

Please hit thumps up if the answer helped you